OpenStudy (anonymous):

Think about a career that interests you in which you would work with very large or very small numbers. Write a paragraph about how you might use scientific notation, standard form, and short word form if you go into that career field. Provide specific examples. Please write at least three complete sentences, using appropriate grammar and punctuation.

OpenStudy (eyust707):

What career are you interested in?

OpenStudy (anonymous):

Um, you should post this in the Writing section. This is math.

Parth (parthkohli):

Not really @julianassange. It's just writing a paragraph on a Math topic.

Parth (parthkohli):

@leondra Are you familiar with scientific notation?

Parth (parthkohli):

Examples: \(100 = 1.0 \times10^2\) \(12818 = 1.2818 \times 10^{4}\) etc.

OpenStudy (anonymous):

Oh. my bad

OpenStudy (eyust707):

\[0.0000000000000000000001 = 1.0 \times 10^{-22}\]

Parth (parthkohli):

Oopsie, though it's not really necessary :)

Parth (parthkohli):

Oh no.

Parth (parthkohli):

Sorry, I confuses myself with some other notation.